# Integrating Dynamical Systems Modeling with Spatiotemporal scRNA-seq Data Analysis

### v1

@@ 3.2.2 Continuous Temporal Dynamics Modeling
where the infimum is taken over all triplets $\left(p,\mathbf{v},g\right)$ such that $p(\cdot,0)=$ $\nu_{0},p(\cdot,1)=\nu_{1},p(\mathbf{x},t)$ absolutely continuous, and
$\partial_{t}p=-\nabla_{\mathbf{x}}\cdot\left(p\mathbf{v}(\mathbf{x},t)\right)+g(\mathbf{x},t)p$ (3.19)
@@
with vanishing boundary condition: $\lim_{|\mathbf{x}|\rightarrow\infty}p(\mathbf{x},t)=0$. Here $\mathbf{v}(\mathbf{x},t)$ is a new vector field, representing the probability flow ODE field.
@@
Thus, the original SDE $\mathrm{d}{\mathbf{x}}_{t}=\left(\mathbf{b}\left({\mathbf{x}}_{t},t\right)\right)\mathrm{d}t+{\sigma}\left(t\right)\mathrm{d}\mathbf{w}_{t}$ now can be transformed into the probability flow ODE
@@
$\mathrm{d}{\mathbf{x}}_{t}=\underbrace{\left(\mathbf{b}\left({\mathbf{x}}_{t},t\right)-\frac{1}{2}\sigma^{2}(t)\nabla_{\mathbf{x}}\log p({\mathbf{x}}_{t},t)\right)}_{\mathbf{v}\left({\mathbf{x}}_{t},t\right)}\mathrm{d}t.$
@@
If the probability flow ODE’s drift $\mathbf{v}(\mathbf{x},t)$, $\sigma(t)$ and the score function $\nabla_{\mathbf{x}}\log p(\mathbf{x},t)$ are specified, then the the drift term $\mathbf{b}(\mathbf{x},t)$ of the SDE can be recovered by $\mathbf{b}(\mathbf{x},t)=\mathbf{v}(\mathbf{x},t)+\frac{1}{2}\sigma^{2}(t)\nabla_{\mathbf{x}}\log p(\mathbf{x},t).$ Therefore, specifying an SDE is equivalent to specifying the probability flow ODE and the score function $\nabla_{\mathbf{x}}\log{p}(\mathbf{x},t)$. One can then use neural networks $\mathbf{v}_{\theta},~{}g_{\theta}$ and $s_{\theta}$ to parameterize $\mathbf{v}(\mathbf{x},t)$, ${g}(\mathbf{x},t)$, and $\frac{1}{2}\sigma^{2}(t)\log{p}(\mathbf{x},t)$ respectively.
@@
To train DeepRUOT, the overall loss is composed of three parts, i.,e., the energy loss, reconstruction loss, and the Fokker-Planck constraint:
$\mathcal{L}=\mathcal{L}_{\text{Energy}}+\lambda_{r}\mathcal{L}_{\text{Recons}}+\lambda_{f}\mathcal{L}_{\text{FP}}.$ (3.20)
@@
The $\mathcal{L}_{\text{Energy}}$ loss aims for the least action of kinetic energy in Equation (3.18), which can be computed via CNF by adopting the similar approach in TIGON (Sha et al., 2024). The reconstruction loss $\mathcal{L}_{\text{Recons}}$ aims the dynamics to match data distribution at the later time point (i.e., $p(\cdot,1)=\nu_{1}$). To achieve the matching in unbalanced settings, DeepRUOT further decomposes it into two parts:
$\mathcal{L}_{\text{Recons}}=\lambda_{m}\mathcal{L}_{\text{Mass}}+\lambda_{d}\mathcal{L}_{\text{OT}}$ (3.21)
@@
where $\mathcal{L}_{\text{Mass}}$ aims to align the number of cells and $\mathcal{L}_{\text{OT}}$ uses normalized weights to perform optimal transport matching. Lastly, $\mathcal{L}_{\text{FP}}$ aims to let the three parameterized neural networks satisfy the Fokker-Planck constraints (3.19). DeepRUOT first utilizes a Gaussian mixture model to estimate the initial distribution, ensuring that it satisfies the initial conditions $p_{0}$, and the physics-informed (PINN) loss is defined as
@@
${\mathcal{L}_{\text{FP}}}=\left\|\partial_{t}p_{\theta}+\nabla_{\mathbf{x}}\cdot\left(p_{\theta}\mathbf{v}_{\theta}\right)-g_{\theta}p_{\theta}\right\|+\lambda_{w}\left\|p_{\theta}(\mathbf{x},0)-p_{0}\right\|,p_{\theta}=\exp{\frac{2}{\sigma^{2}}s_{\theta}}.$ (3.22)
@@
In (Zhang et al., 2025), DeepRUOT adopts a two-stage training approach to stabilize the training process. For the pre-training stage, they use reconstruction loss only to train $\mathbf{v}_{\theta}$ and $g_{\theta}$. Then, they fix $\mathbf{v}_{\theta}$ and $g_{\theta}$ and employ conditional flow-matching (Lipman et al., 2023; Tong et al., 2024a, b) to learn the log density function ($s_{\theta}(\mathbf{x},t)$). Finally, for the training stage, they use the $\mathbf{v}_{\theta}$, $g_{\theta}$, and the log density function as the starting point, then obtain the final result by minimizing the total loss (3.20).
@@ 4 Dynamic Modeling of Spatial Transcriptomics
In this section, we review the dynamical modeling approaches for spatial transcriptomics data. We will first present several random walk or ODE-based methods to model the snapshot spatial data. Next, we focus on the recent progress to dissect the spatio-temporal dynamics underlying datasets with both space and time resolutions.
@@ 4.1 Snapshot Spatial Transcriptomics
Below we describe several modeling strategies for single snapshot spatial transcriptomics data, including pseudotime, random walk, and continuous differential models respectively.
@@ 4.1.1 Pseudotime Methods
In the context of spatiotemporal trajectory inference, stLearn (Pham et al., 2023) proposes the PSTS algorithm that combines spatial information and geodesic-based pseudotime information to infer the spatiotemporal developmental trajectory of cells. The pseudotime distance between two clusters is defined as
@@ 4.1.2 Discrete Spatial Dynamics Modeling
STT (Zhou et al., 2024c) is a random-walk-based algorithm to detect multi-stable attractors in spatial transcriptomics. Central to STT the incorporation of a space coordinate-aware random walk, with the transition probability matrix has the form $P=w_{1}P_{v}+w_{2}P_{c}+(1-w_{1}-w_{2})P_{s}$, where $P_{v}$ is induced by an attractor-specific RNA velocity (named spatial transition tensor), $P_{c}$ is induced by gene expression similarity (i.e. diffusion in the gene space), and $P_{s}$ is induced by space coordinates (i.e. diffusion in the physical space). Iteratively 1) decomposing $P$ to identify attractors and assign attractor membership to each individual cell and 2) improving attractor-specific RNA velocity estimation, STT is able to identify transitional cells in the snapshot spatial transcriptomics data and plot the local streamlines within attractors.
SpaTrack (Shen et al., 2025) is a spatial transcriptomics analysis tool based on optimal transport theory, which reconstructs cell differentiation trajectories by integrating gene expression profiles and spatial coordinates of cells. When processing Snapshot data, SpaTrack defines the transition cost matrix between cells by weighting gene expression distance and spatial distance $C_{ij}=\alpha_{1}\|\mathbf{g}_{i}-\mathbf{g}_{j}\|^{2}+\alpha_{2}\|\mathbf{z}_{i}-\mathbf{z}_{j}\|^{2}$ where $\mathbf{g}_{i}$ represents the gene expression of cell $i$, $\mathbf{x}_{i}$ denotes the spatial coordinates of cell $i$, and $\alpha_{1},\alpha_{2}$ are weighting coefficients. The transition probability matrix between cells is obtained by solving the following entropy-regularized optimal transport (OT) problem
$\mathbf{P}=\arg\min_{\mathbf{P}}\sum_{ij}C_{ij}P_{ij}+\epsilon H(\mathbf{P})\quad\mathrm{s.t.}\sum_{i}P_{ij}=1,\sum_{j}P_{ij}=1$
where $H(\mathbf{P})$ denotes the entropy regularization term and $\epsilon$ is the regularization coefficient. SpaTrack identifies trajectory starting points using single-cell entropy. Let the identified starting points be cells $1,2,\cdots,s$. The probability of transitioning from starting cells to cell $i$ can be calculated as $\gamma_{i}=\sum_{j=1}^{s}P_{ji}$. By sorting cells in ascending order based on their $\gamma_{i}$ values, the position of each cell in the differentiation trajectory can be determined.
@@ 4.1.3 Continuous Spatial Dynamics Modeling
Several methods aim to extend the continuous RNA velocity model of scRNA-seq data toward snapshot spatial transcriptomics. One recent method, iSORT (Tan et al., 2024) uses transfer learning to obtain a mapping of gene expression to spatial location $\mathbf{z}=f(\mathbf{x})$ and proposes the concept of spatial RNA velocity, which utilizes the velocity of gene expression and the mapping $\mathbf{z}=f(\mathbf{x})$ to obtain spatial RNA velocity, formally
$\frac{\mathrm{d}\mathbf{z}}{\mathrm{d}t}=\nabla f\cdot\frac{\mathrm{d}\mathbf{x}}{\mathrm{d}t}.$
In addition, Topovelo (Gu et al., 2024b) uses a graph neural network to infer RNA velocity for spatial transcriptomics data and suggested that a decoder could be trained to further infer continuous spatial velocities.
@@ 4.2 Temporally Resolved Spatial Transcriptomics
The availability of time-series ST data opens new avenues to explore cellular migration within physical space (Peng et al., 2024; Qiu et al., 2022b; Gu et al., 2024b). Nevertheless, the inherently destructive nature of sequencing limits ST data to static snapshots rather than continuous trajectories. Particularly, when sequencing is performed at various time points during embryonic development, the resulting time-series ST data are often derived from distinct biological samples, leading to multiple unpaired snapshots (Chen et al., 2022a; Wei et al., 2022; Wang et al., 2024b). In addition, due to possible rotation, translation, and stretching of different slices, the spatial coordinates of different samples are not in the same coordinate system (Zeira et al., 2022; Qiu et al., 2022b; Peng et al., 2024). Therefore, reconstructing trajectories of cell state transition, proliferation, and migration for time-series ST data is a challenging task.
To overcome these challenges, many methods have been developed in recent years. Similar to modeling temporal single-cell data, these methods can be divided into two categories: those that model dynamics on discrete cell states (Klein et al., 2025; Halmos et al., 2024; Qiu et al., 2022b) and those that model dynamics in continuous spaces (Peng et al., 2024).
@@ 4.2.1 Discrete Spatiotemporal Dynamics Modeling
Among the methods that model spatiotemporal dynamics on discrete cell states, recent work includes Moscot (Klein et al., 2025), DeST-OT (Halmos et al., 2024) and Spateo (Qiu et al., 2022b). These approaches employ fused Gromov-Wasserstein optimal transport (Titouan et al., 2019; Chowdhury and Mémoli, 2019) as the main tool, which was first used by PASTE (Zeira et al., 2022) to align adjacent 2D slices to reconstruct the 3D structure of the tissue.
Fused Gromov-Wasserstein Optimal Transport
We consider two adjacent unpaired slices $(\mathbf{X},\mathbf{Z})$ and $(\mathbf{X}^{\prime},\mathbf{Z}^{\prime})$ with spots (or cell) numbers $N$ and $M$, where $\mathbf{X}\in\mathbb{R}^{N\times G}$, $\mathbf{X}^{\prime}\in\mathbb{R}^{M\times G}$ are the gene expression of the two slices, and $\mathbf{Z}\in\mathbb{R}^{N\times 2}$ and $\mathbf{Z}^{\prime}\in\mathbb{R}^{M\times 2}$ are the spatial coordinates. In addition, the spatial coordinates of each slice can be converted into a distance matrix $\mathbf{D}\in\mathbb{R}^{N\times N}_{+}$, where $d_{ij}=\|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}$. The fused Gromov-Wasserstein optimal transport problem reads:
$\min_{\boldsymbol{\pi}\in\boldsymbol{\pi}(\nu_{0},\nu_{1})}(1-\alpha)\sum_{i,j}c_{i,j}\pi_{i,j}+\alpha\sum_{i,j,k,l}(d_{i,k}-d_{j,l}^{\prime})^{2}\pi_{i,j}\pi_{k,l},$ (4.23)
where $\pi,\boldsymbol{\pi}(\nu_{0},\nu_{1})$ and $c$ have the same meaning as before, and $\alpha$ is a hyperparameter that weighs the importance of gene expression and spatial location. The fused Gromov-Wasserstein optimal transport (FGW-OT) problems can also be solved by calling the POT package (Flamary et al., 2021).
Generalized Weighted Procrustes Problem
When the mapping $\boldsymbol{\pi}$ is found, in order to unify the spatial coordinates of adjacent slices into the same coordinate system by a rigid body transformation (rotation and translation), we need to solve a generalized weighted Procrustes problem. Formally,
$\hat{\mathbf{R}},\hat{\mathbf{r}}=\mathop{\arg\min}_{\begin{subarray}{c}\mathbf{R}\in\mathbb{R}^{2\times 2},\mathbf{r}\in\mathbb{R}^{2}\\ \mathbf{R}^{T}\mathbf{R}=\mathbf{I},\det\mathbf{R}=1\end{subarray}}\sum_{i,j}\pi_{ij}\left\|\mathbf{z}_{i}-(\mathbf{R}\mathbf{z}^{\prime}_{j}+\mathbf{r})\right\|^{2}_{2},$ (4.24)
where $\mathbf{R}$ is the rotation matrix and $\mathbf{r}$ is the translation vector.
Applications in ST Data
Moscot (Klein et al., 2025) extends FGW-OT to model slices of adjacent time points by adding the entropy regularization mentioned in Eq. (3.7) and the unbalanced settings mentioned in Eqs. (3.8) and (3.9). Formally,
@@
$+\tau_{1}\operatorname{KL}(\boldsymbol{\pi}\mathbf{1}_{M}||\nu_{0})+\tau_{2}\operatorname{KL}(\boldsymbol{\pi}^{T}\mathbf{1}_{N}||\nu_{1})-\varepsilon H(\boldsymbol{\pi}),$
@@
where $\tau_{1},\tau_{2},\varepsilon$ and $H(\boldsymbol{\pi})$ have the same meaning as before and $\nu_{0}$ is calculated from a pre-selected gene set according to Eq. (3.9). Note that when we talk about ST data at different time points, the spatial coordinates can not only be 2D but can also be reconstructed in 3D. We use $d_{\text{spa}}$ to refer to the dimension of the spatial coordinates.
@@
DeST-OT (Halmos et al., 2024) designs methods that enable simultaneous inference of cell growth rate and mapping from data. The DeST-OT optimization problem with the semi-relaxed constraints and entropic regularization is
@@
$\begin{array}[]{ll}\min&\mathcal{E}_{\text{DeST-OT }}(\boldsymbol{\pi})+\tau_{1}\operatorname{KL}(\boldsymbol{\pi}\mathbf{1}_{M}||\nu_{0})-\varepsilon H(\boldsymbol{\pi})\\ \text{ s.t. }&\boldsymbol{\pi}^{T}\mathbf{1}_{N}=\nu_{1},\quad\boldsymbol{\pi}\in\mathbb{R}^{N\times M}_{+},\end{array}$
@@
where $\mathcal{E}_{\text{DeST-OT }}(\boldsymbol{\pi})$ includes the Wasserstein-OT term and another term $\mathcal{E}^{\mathbf{M}}(\boldsymbol{\pi})$ related to growth and GW-OT, that is
$\mathcal{E}_{\text{DeST-OT }}:=(1-\alpha)\sum_{i,j^{\prime}}c_{i,j^{\prime}}\pi_{i,j^{\prime}}+\alpha\mathcal{E}^{\mathbf{M}}(\boldsymbol{\pi}).$
$\mathcal{E}^{\mathbf{M}}(\boldsymbol{\pi})$ is defined as
$\mathcal{E}^{\mathbf{M}}(\boldsymbol{\pi}):=\frac{1}{2}$ $\left(\sum_{i,j^{\prime},k^{\prime}}\pi_{ij^{\prime}}\pi_{ik^{\prime}}\mathbf{M}_{j^{\prime}k^{\prime}}^{\prime 2}+\sum_{i,j,k^{\prime}}\pi_{ik^{\prime}}\pi_{jk^{\prime}}\mathbf{M}_{ij}^{2}\right.$
$\left.+\sum_{i,j^{\prime},k,l^{\prime}}\left(\mathbf{M}_{ik}-\mathbf{M}_{j^{\prime}l^{\prime}}^{\prime}\right)^{2}\pi_{ij^{\prime}}\pi_{kl^{\prime}}\right),$

### v2

@@ 3.2.2 Continuous Temporal Dynamics Modeling
where the infimum is taken over all triplets $\left(p,\boldsymbol{v},g\right)$ such that $p(\cdot,0)=$ $\nu_{0},p(\cdot,1)=\nu_{1},p(\boldsymbol{x},t)$ absolutely continuous, and
$\partial_{t}p=-\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{v}(\boldsymbol{x},t)\right)+g(\boldsymbol{x},t)p$ (20)
@@
with vanishing boundary condition: $\lim_{|\boldsymbol{x}|\rightarrow\infty}p(\boldsymbol{x},t)=0$. Here $\boldsymbol{v}(\boldsymbol{x},t)$ is a new vector field, representing the probability flow ODE field.
@@
Thus, the original SDE $\mathrm{d}{\boldsymbol{x}}_{t}=\left(\boldsymbol{b}\left({\boldsymbol{x}}_{t},t\right)\right)\mathrm{d}t+{\sigma}\left(t\right)\mathrm{d}\boldsymbol{w}_{t}$ now can be transformed into the probability flow ODE
@@
$\mathrm{d}{\boldsymbol{x}}_{t}=\underbrace{\left(\boldsymbol{b}\left({\boldsymbol{x}}_{t},t\right)-\frac{1}{2}\sigma^{2}(t)\nabla_{\boldsymbol{x}}\log p({\boldsymbol{x}}_{t},t)\right)}_{\boldsymbol{v}\left({\boldsymbol{x}}_{t},t\right)}\mathrm{d}t.$
@@
If the probability flow ODE’s drift $\boldsymbol{v}(\boldsymbol{x},t)$, $\sigma(t)$ and the score function $\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x},t)$ are specified, then the the drift term $\boldsymbol{b}(\boldsymbol{x},t)$ of the SDE can be recovered by $\boldsymbol{b}(\boldsymbol{x},t)=\boldsymbol{v}(\boldsymbol{x},t)+\frac{1}{2}\sigma^{2}(t)\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x},t).$ Therefore, specifying an SDE is equivalent to specifying the probability flow ODE and the score function $\nabla_{\boldsymbol{x}}\log{p}(\boldsymbol{x},t)$. One can then use neural networks $\boldsymbol{v}_{\theta},~{}g_{\theta}$ and $s_{\theta}$ to parameterize $\boldsymbol{v}(\boldsymbol{x},t)$, ${g}(\boldsymbol{x},t)$, and $\frac{1}{2}\sigma^{2}(t)\log{p}(\boldsymbol{x},t)$, respectively.
@@
To train DeepRUOT, the overall loss is composed of three parts, i.e., the energy loss, reconstruction loss, and the Fokker–Planck constraint:
$\mathcal{L}=\mathcal{L}_{\text{Energy}}+\lambda_{r}\mathcal{L}_{\text{Recons}}+\lambda_{f}\mathcal{L}_{\text{FP}}.$ (21)
@@
The $\mathcal{L}_{\text{Energy}}$ loss aims for the least action of kinetic energy in Equation (19), which can be computed via CNF by adopting the similar approach in TIGON Sha et al. (2024). The reconstruction loss $\mathcal{L}_{\text{Recons}}$ aims the dynamics to match data distribution at the later time point (i.e., $p(\cdot,1)=\nu_{1}$). To achieve the matching in unbalanced settings, DeepRUOT further decomposes it into two parts:
$\mathcal{L}_{\text{Recons}}=\lambda_{m}\mathcal{L}_{\text{Mass}}+\lambda_{d}\mathcal{L}_{\text{OT}}$ (22)
@@
where $\mathcal{L}_{\text{Mass}}$ aims to align the number of cells and $\mathcal{L}_{\text{OT}}$ uses normalized weights to perform optimal transport matching. Lastly, $\mathcal{L}_{\text{FP}}$ aims to let the three parameterized neural networks satisfy the Fokker–Planck constraints (20). DeepRUOT first utilizes a Gaussian mixture model to estimate the initial distribution, ensuring that it satisfies the initial conditions $p_{0}$, and the physics-informed (PINN) loss Raissi et al. (2019) is defined as
@@
${\mathcal{L}_{\text{FP}}}=\left\|\partial_{t}p_{\theta}+\nabla_{\boldsymbol{x}}\cdot\left(p_{\theta}\boldsymbol{v}_{\theta}\right)-g_{\theta}p_{\theta}\right\|+\lambda_{w}\left\|p_{\theta}(\boldsymbol{x},0)-p_{0}\right\|,p_{\theta}=\exp{\frac{2}{\sigma^{2}}s_{\theta}}.$ (23)
@@
In Zhang et al. (2025), DeepRUOT adopts a two-stage training approach to stabilize the training process. For the pre-training stage, they use reconstruction loss only to train $\boldsymbol{v}_{\theta}$ and $g_{\theta}$. Then, they fix $\boldsymbol{v}_{\theta}$ and $g_{\theta}$ and employ conditional flow-matching Lipman et al. (2023); Tong et al. (2024, 2024) to learn the log density function ($s_{\theta}(\boldsymbol{x},t)$). Finally, for the training stage, they use the $\boldsymbol{v}_{\theta}$, $g_{\theta}$, and the log density function as the starting point, then obtain the final result by minimizing the total loss (21).
@@ 4 Dynamic Modeling of Spatial Transcriptomics
In this section, we review the dynamical modeling approaches for spatial transcriptomics data. We will first present several random walk- or ODE-based methods to model the snapshot spatial data. Next, we focus on the recent progress to dissect the spatiotemporal dynamics underlying datasets with both space and time resolutions. Figure 4 provides an overview of these spatial transcriptomics modeling approaches.
{adjustwidth}
[…]
@@ 4.1 Snapshot Spatial Transcriptomics
Below we describe several modeling strategies for single snapshot spatial transcriptomics data, including pseudotime, random walk, and continuous differential equation models, respectively.
@@ 4.1.1 Pseudotime Methods
In the context of spatiotemporal trajectory inference, stLearn Pham et al. (2023) proposes the PSTS algorithm that combines spatial information and geodesic-based pseudotime information to infer the spatiotemporal developmental trajectory of cells. The pseudotime distance between two clusters is defined as
@@ 4.1.2 Discrete Spatial Dynamics Modeling
STT Zhou et al. (2024) is a random walk-based algorithm to detect multi-stable attractors in spatial transcriptomics. Central to STT is the incorporation of a space coordinate-aware random walk, with the transition probability matrix having the form $P=w_{1}P_{v}+w_{2}P_{c}+(1-w_{1}-w_{2})P_{s}$, where $P_{v}$ is induced by an attractor-specific RNA velocity (named spatial transition tensor), $P_{c}$ is induced by gene expression similarity (i.e., diffusion in the gene space), and $P_{s}$ is induced by space coordinates (i.e., diffusion in the physical space). By iteratively (1) decomposing $P$ to identify attractors and assign attractor membership to each individual cell and (2) improving attractor-specific RNA velocity estimation, STT is able to identify transitional cells in the snapshot spatial transcriptomics data and plot the local streamlines within attractors.
SpaTrack Shen et al. (2025) is a spatial transcriptomics analysis tool based on optimal transport theory, which reconstructs cell differentiation trajectories by integrating gene expression profiles and spatial coordinates of cells. When processing Snapshot data, SpaTrack defines the transition cost matrix between cells by weighting gene expression distance and spatial distance as follows: $C_{ij}=\alpha_{1}\|\boldsymbol{g}_{i}-\boldsymbol{g}_{j}\|^{2}+\alpha_{2}\|\boldsymbol{z}_{i}-\boldsymbol{z}_{j}\|^{2}$ where $\boldsymbol{g}_{i}$ represents the gene expression of cell $i$, $\boldsymbol{x}_{i}$ denotes the spatial coordinates of cell $i$, and $\alpha_{1},\alpha_{2}$ are weighting coefficients. The transition probability matrix between cells is obtained by solving the following entropy-regularized optimal transport (OT) problem
$\boldsymbol{P}=\arg\min_{\boldsymbol{P}}\sum_{ij}C_{ij}P_{ij}+\epsilon H(\boldsymbol{P})\quad\mathrm{s.t.}\sum_{i}P_{ij}=1,\sum_{j}P_{ij}=1$
where $H(\boldsymbol{P})$ denotes the entropy regularization term and $\epsilon$ is the regularization coefficient. SpaTrack identifies trajectory starting points using single-cell entropy. Let the identified starting points be cells $1,2,\cdots,s$. The probability of transitioning from starting cells to cell $i$ can be calculated as $\gamma_{i}=\sum_{j=1}^{s}P_{ji}$. By sorting cells in ascending order based on their $\gamma_{i}$ values, the position of each cell in the differentiation trajectory can be determined.
@@ 4.1.3 Continuous Spatial Dynamics Modeling
Several methods aim to extend the continuous RNA velocity model of scRNA-seq data toward snapshot spatial transcriptomics. One recent method, iSORT  Tan et al. (2024) uses transfer learning to obtain a mapping of gene expression to spatial location $\boldsymbol{z}=f(\boldsymbol{x})$ and proposes the concept of spatial RNA velocity, which utilizes the velocity of gene expression and the mapping $\boldsymbol{z}=f(\boldsymbol{x})$ to obtain spatial RNA velocity, formally
$\frac{\mathrm{d}\boldsymbol{z}}{\mathrm{d}t}=\nabla f\cdot\frac{\mathrm{d}\boldsymbol{x}}{\mathrm{d}t}.$
In addition, Topovelo  Gu et al. (2024) uses a graph neural network to infer RNA velocity for spatial transcriptomics data and suggests that a decoder could be trained to further infer continuous spatial velocities.
@@ 4.2 Temporally Resolved Spatial Transcriptomics
The availability of time-series ST data opens new avenues to explore cellular migration within physical space Peng et al. (2024); Qiu et al. (2022); Gu et al. (2024). Nevertheless, the inherently destructive nature of sequencing limits ST data to static snapshots rather than continuous trajectories. Particularly, when sequencing is performed at various time points during embryonic development, the resulting time-series ST data are often derived from distinct biological samples, leading to multiple unpaired snapshots Chen et al. (2022); Wei et al. (2022); Wang et al. (2024). In addition, due to possible rotation, translation, and stretching of different slices, the spatial coordinates of different samples are not in the same coordinate system Zeira et al. (2022); Qiu et al. (2022); Peng et al. (2024). Therefore, reconstructing trajectories of cell state transition, proliferation, and migration for time-series ST data is a challenging task.
To overcome these challenges, many methods have been developed in recent years. Similar to modeling temporal single-cell data, these methods can be divided into two categories: those that model dynamics on discrete cell states Klein et al. (2025); Halmos et al. (2024); Qiu et al. (2022) and those that model dynamics in continuous spaces Peng et al. (2024).
@@ 4.2.1 Discrete Spatiotemporal Dynamics Modeling
Among the methods that model spatiotemporal dynamics on discrete cell states, recent work includes Moscot Klein et al. (2025), DeST-OT Halmos et al. (2024), and Spateo Qiu et al. (2022). These approaches employ fused Gromov–Wasserstein optimal transport Titouan et al. (2019); Chowdhury and Mémoli (2019) as the main tool, which was first used by PASTE Zeira et al. (2022) to align adjacent 2D slices to reconstruct the 3D structure of the tissue.
Fused Gromov–Wasserstein Optimal Transport We consider two adjacent unpaired slices $(\boldsymbol{X},\boldsymbol{Z})$ and $(\boldsymbol{X}^{\prime},\boldsymbol{Z}^{\prime})$ with spots (or cell) numbers $N$ and $M$, where $\boldsymbol{X}\in\mathbb{R}^{N\times G}$, $\boldsymbol{X}^{\prime}\in\mathbb{R}^{M\times G}$ are the gene expression of the two slices, and $\boldsymbol{Z}\in\mathbb{R}^{N\times 2}$ and $\boldsymbol{Z}^{\prime}\in\mathbb{R}^{M\times 2}$ are the spatial coordinates. In addition, the spatial coordinates of each slice can be converted into a distance matrix $\boldsymbol{D}\in\mathbb{R}^{N\times N}_{+}$, where $d_{ij}=\|\boldsymbol{z}_{i}-\boldsymbol{z}_{j}\|_{2}$. The fused Gromov–Wasserstein optimal transport problem reads as follows:
$\min_{\boldsymbol{\pi}\in\boldsymbol{\Pi}(\nu_{0},\nu_{1})}(1-\alpha)\sum_{i,j}c_{i,j}\pi_{i,j}+\alpha\sum_{i,j,k,l}(d_{i,k}-d_{j,l}^{\prime})^{2}\pi_{i,j}\pi_{k,l},$ (24)
where $\pi,\boldsymbol{\Pi}(\nu_{0},\nu_{1})$ and $c$ have the same meaning as before, and $\alpha$ is a hyperparameter that weighs the importance of gene expression and spatial location. The fused Gromov–Wasserstein optimal transport (FGW-OT) problems can also be solved by calling the POT (version 0.9.5) package Flamary et al. (2021).
Generalized Weighted Procrustes Problem When the mapping $\boldsymbol{\pi}$ is found, in order to unify the spatial coordinates of adjacent slices into the same coordinate system by a rigid body transformation (rotation and translation), we need to solve a generalized weighted Procrustes problem. Formally,
$\hat{\boldsymbol{R}},\hat{\boldsymbol{r}}=\mathop{\arg\min}_{\begin{subarray}{c}\boldsymbol{R}\in\mathbb{R}^{2\times 2},\boldsymbol{r}\in\mathbb{R}^{2}\\ \boldsymbol{R}^{T}\boldsymbol{R}=\boldsymbol{I},\det\boldsymbol{R}=1\end{subarray}}\sum_{i,j}\pi_{ij}\left\|\boldsymbol{z}_{i}-(\boldsymbol{R}\boldsymbol{z}^{\prime}_{j}+\boldsymbol{r})\right\|^{2}_{2},$ (25)
where $\boldsymbol{R}$ is the rotation matrix and $\boldsymbol{r}$ is the translation vector.
Applications in ST Data Moscot Klein et al. (2025) extends FGW-OT to model slices of adjacent time points by adding the entropy regularization mentioned in Equation (8) and the unbalanced settings mentioned in Equations (9) and (10). Formally,
@@
$+\tau_{1}\operatorname{KL}(\boldsymbol{\pi}\boldsymbol{1}_{M}||\nu_{0})+\tau_{2}\operatorname{KL}(\boldsymbol{\pi}^{T}\boldsymbol{1}_{N}||\nu_{1})-\varepsilon H(\boldsymbol{\pi}),$
@@
where $\tau_{1},\tau_{2},\varepsilon$ and $H(\boldsymbol{\pi})$ have the same meaning as before and $\nu_{0}$ is calculated from a pre-selected gene set according to Equation (10). Note that when we talk about ST data at different time points, the spatial coordinates can not only be 2D but can also be reconstructed in 3D. We use $d_{\text{spa}}$ to refer to the dimension of the spatial coordinates.
@@
DeST-OT Halmos et al. (2024) designs methods that enable simultaneous inference of cell growth rate and mapping from data. The DeST-OT optimization problem with the semi-relaxed constraints and entropic regularization is
@@
$\begin{array}[]{ll}\min&\mathcal{E}_{\text{DeST-OT }}(\boldsymbol{\pi})+\tau_{1}\operatorname{KL}(\boldsymbol{\pi}\boldsymbol{1}_{M}||\nu_{0})-\varepsilon H(\boldsymbol{\pi})\\ \text{ s.t. }&\boldsymbol{\pi}^{T}\boldsymbol{1}_{N}=\nu_{1},\quad\boldsymbol{\pi}\in\mathbb{R}^{N\times M}_{+},\end{array}$
@@
where $\mathcal{E}_{\text{DeST-OT }}(\boldsymbol{\pi})$ includes the Wasserstein OT term and another term $\mathcal{E}^{\boldsymbol{M}}(\boldsymbol{\pi})$ related to growth and GW-OT, that is,
$\mathcal{E}_{\text{DeST-OT }}:=(1-\alpha)\sum_{i,j^{\prime}}c_{i,j^{\prime}}\pi_{i,j^{\prime}}+\alpha\mathcal{E}^{\boldsymbol{M}}(\boldsymbol{\pi}).$
$\mathcal{E}^{\boldsymbol{M}}(\boldsymbol{\pi})$ is defined as
$\mathcal{E}^{\boldsymbol{M}}(\boldsymbol{\pi}):=\frac{1}{2}$ $\left(\sum_{i,j^{\prime},k^{\prime}}\pi_{ij^{\prime}}\pi_{ik^{\prime}}\boldsymbol{M}_{j^{\prime}k^{\prime}}^{\prime 2}+\sum_{i,j,k^{\prime}}\pi_{ik^{\prime}}\pi_{jk^{\prime}}\boldsymbol{M}_{ij}^{2}\right.$
$\left.+\sum_{i,j^{\prime},k,l^{\prime}}\left(\boldsymbol{M}_{ik}-\boldsymbol{M}_{j^{\prime}l^{\prime}}^{\prime}\right)^{2}\pi_{ij^{\prime}}\pi_{kl^{\prime}}\right),$